# Efficient Response Generation Strategy Selection for Fine-Tuning Large Language Models Through Self-Aligned Perplexity

### v2

@@ 5 Experiment
In this section, we treat each generation strategy from Section 4.3, along with other response-selection methods mentioned in the related work section, as baselines. We then compare the average training outcomes of our method against these baselines across all tasks.
When training the target model using a particular data generation strategy, we run experiments with random seeds 0, 1, and 2, and report their average performance in Table 9. For our method, we also repeat the process three times, each time selecting a different subset of size from the training data, and then report the average of these three runs. For instance, one run may use the first 50 samples, another the second 50, and so on, and we combine their results to obtain a final average.
@@ 5.1 Hyperparameters
5.2 Our Method vs. Baselines (Each Data Generation Methods)
Figure 1 compares our method against baselines. The Rewrite Ground Truth method is omitted from the Figure 1 and Figure 2 since it applies only to datasets with CoT groundtruth. Other results represent task averages across all datasets, while Rewrite Ground Truth averages a subset of tasks. Despite its exclusion from the figure, all metrics and tables incorporate this method’s results.
As shown in Figure 1, our method outperforms every data generation strategy. On average, it surpasses the best data generation strategy (Claude) by 0.66%. The average performance of all methods is 62.97%, which is 1.22% lower than ours (64.19%).
5.3 Choosing Data Generation Methods Becomes More Important When PGR Increases
When analyzing the detailed performance record (n = 1000) (Table 9 in Appendix), we notice that there is substantial performance variation between methods for some tasks, while for others, the variation is minimal. For example, when training the Qwen model on API-BANK, using training data generated by Claud Answer Directly strategy results in an evaluation accuracy of 45.7%, whereas training with data generated by GPT-4o Examples strategy yields an evaluation accuracy of 33.9%. In contrast, for tasks like GSM8K and MATH Algebra, the performance gap between the best and worst methods is under 2%.
We observe that the performance variation between methods is minimal when training with 1000 samples (data from Table 9 in Appendix) yields only negligible improvement over training with 100 samples (data from Table 13 in Appendix).
To investigate this relationship, we created Figure 3, where the X-axis represents the performance gain ratio (PGR) on a logarithmic scale, and the Y-axis measures the standard diviation in training effectiveness across methods. The PGR is defined as:
This metric quantifies the improvement in accuracy when increasing the training data from 100 to 1000 examples.
While the right plot demonstrates that when the PGR exceeds 2%, the diversity across generation methods increases significantly, the left plot shows a strong positive correlation between and . Notably, when PGR is below 2%, the diversity is nearly 0.
These two subplots show that the variation in training effectiveness increases as PGR increases. The greater the divergence between the training outcomes, the more important it becomes to select the best response generation strategy.
5.3.1 Our Method Performs Better as PGR Increases
We investigate whether our method becomes more effective as the dataset is filtered using increasingly higher PGR thresholds. Figure 2 compares our method with baselines on subsets where the minimum PGR ranges from (all data) to 0.3. As the PGR threshold increases, the performance gap between our method and the average baseline grows from 1.3% to 3.5%, indicating that our method performs especially well on datasets with higher PGR values.
5.3.2 Comparison with Other Data Selection Metrics
Figure 4 shows that our method consistently outperform other methods in terms of accuracy and weighted spearman correlation. As the PGR threshold increases, this gap widens.
@@ 6.1 Why Self-Aligned Perplexity Outperforms Traditional Perplexity
| Target Response style | Model | Task | PPL | ICPPL | ICPPL | ICPPL |
@@
Failure example of perplexity: As shown in Table 2, on the ECQA and PIQA tasks, responses written in a highly redundant style (see Appendix A.7) receive lower perplexity scores—even though the target model is unlikely to generate such responses—while high-quality GPT-4o Answer Directly responses receive higher perplexity.
Here’s a slightly more concise version with one word removed for brevity:
Effectiveness of in-context example: When we incorporate a style-specific in-context example (e.g., a step-by-step response), in-context perplexity adjusts the model’s probability estimation to favor responses that match the provided style. Our method uses the target model’s own prediction as the in-context example, allowing self-aligned perplexity to better reflect the alignment between the model’s style and the response candidates.
@@
| Ours | Mistral7B | 0.582 | 0.411 |
| Train-then-test lr = 2e-5 | 0.578 | 0.249 | |
| Train-then-test lr = 2e-4 | 0.579 | 0.440 | |
| Ours | Llama3 | 0.650 | 0.344 |
| Train-then-test lr = 2e-5 | 0.626 | 0.257 | |
| Train-then-test lr = 2e-4 | 0.638 | 0.286 | |
| Ours | Qwen2.5 | 0.723 | 0.494 |
| Train-then-test lr = 2e-5 | 0.710 | 0.160 | |
| Train-then-test lr = 2e-4 | 0.726 | 0.345 | |
| Ours | Average | 0.652 | 0.416 |
| Train-then-test lr = 2e-5 | 0.638 | 0.222 | |
| Train-then-test lr = 2e-4 | 0.648 | 0.357 |
| Model | Accuracy | Weighted Spearman’s | ||||||||
| 0–10 | 10–20 | 20–30 | 0–50 | 0–100 | 0–10 | 10–20 | 20–30 | 0–50 | 0–100 | |
| Mistral7B | 0.583 | 0.584 | 0.570 | 0.583 | 0.583 | 0.396 | 0.310 | 0.257 | 0.412 | 0.411 |
| Llama3 | 0.644 | 0.650 | 0.642 | 0.650 | 0.652 | 0.331 | 0.393 | 0.077 | 0.344 | 0.314 |
| Qwen2.5 | 0.714 | 0.711 | 0.718 | 0.723 | 0.726 | 0.363 | 0.472 | 0.327 | 0.494 | 0.401 |
@@
One natural—but computationally expensive—approach to selecting the optimal response generation strategy is to adopt a ”Train-Then-Select” procedure. In this approach, we first generate a small dataset (e.g., 100 samples) using each candidate strategy. For each dataset, we train the target model and evaluate its performance. We then rank the strategies based on these results and choose the best-performing one to generate the remainder of the dataset.
When evaluating ”Train-Then-Select”, we train the target model on 100 samples under two settings:
1. Standard Training: A learning rate of 2e-5 for 20 epochs (matching our main setup). The performance accuracies for each strategy under this settings is reported in the Table 13.
2. Intense Training: A learning rate of 2e-4 for 40 epochs. The performance accuracies for each strategy under this settings is reported in the Table 14.
After ranking the strategies using Train-Then-Select, we compare its performance with our method. As shown in Table 3, despite using less data and requiring no training, validation, or testing computations for strategy selection, our method achieves better accuracy and higher Spearman correlation.
6.3 Stability of Our Methods
As shown in Table 4, our method exhibits improved accuracy as the subset size increases. Generally speaking, the performance of our method is consistent. Some of the small subsets might causes degradtion of the accuracy(highlighted in red), thus we recommend to choose a relatively larger size of subsets. In the table, ”0–10”, ”10–20”, and ”20–30” indicate that we use the first, second, and third groups of 10 training samples, respectively, to compute the ranking. ”0–50” and ”0–100” represent subsets containing the first 50 and 100 samples. According to the table, when the subset is large (e.g., K = 50 or K = 100), the performance is better in terms of accuracy and Spearman correlation. However, when the subset is small (K = 10), the performance of each subset becomes less stable.
6.4 Groundtruth vs. Synthetic Data
As shown in Table 9 of Appendix, when ground truth is provided in natural language (e.g., GSM8K, MATH, ECQA, MBPP), training on ground truth is less effective than training on synthetic data. This is because LLMs are more familiar with LLM-generated data, as demonstrated by Ren et al. (2024). However, when the ground truth is written as a gold label without a CoT inference process, training on gold label can sometimes outperform training on CoT synthetic data within the same domain. However, as shown in Table 7 of Appendix, training on gold labels harms cross-domain performance more than training on synthetic data. Additionally, in real-life scenarios, training on natural language data is crucial, as users expect to see the rationale behind the final prediction made by LLMs.

### v3

@@ 5 Experiment
| Methods | STD Range | num of recorded data | mistral | llama 3 instruct | qwen | Avg Acc |
| Upper bound | All Data | 51.0 | 59.39% | 64.44% | 71.41% | 65.08% |
| Step-by-step | 56.25% | 60.94% | 69.87% | 62.35% | ||
| GPT-4 ICL examples | 57.29% | 62.03% | 69.90% | 63.07% | ||
| Human examples | 56.95% | 61.91% | 70.22% | 63.03% | ||
| Mini-GPT-4 | 56.72% | 61.36% | 70.13% | 62.74% | ||
| GPT-4 | 56.92% | 62.83% | 69.89% | 63.21% | ||
| Claude | 57.45% | 62.93% | 70.30% | 63.56% | ||
| Ours | 58.34% | 63.66% | 70.50% | 64.16% | ||
| Ours - Claude | All Data | 51.0 | +0.88% | +0.72% | +0.19% | +0.60% |
| Ours - Avg of Others | +1.41% | +1.66% | +0.45% | +1.17% | ||
| Ours - Claude | 21.0 | +1.67% | +1.14% | -0.21% | +1.18% | |
| Ours - Avg of Others | +2.17% | +2.84% | +1.88% | +2.41% | ||
| Ours - Claude | 10.0 | +3.61% | +1.56% | -1.63% | +2.26% | |
| Ours - Avg of Others | +3.60% | +6.00% | +3.38% | +4.54% |
| Methods | STD Range | num of recorded data | mistral | llama 3 instruct | qwen | Avg Acc | Weighted Spearman Pho |
| Upper bound | All Data | 51.0 | 59.39% | 64.44% | 71.41% | 65.08% | |
| skywork | 56.32% | 61.64% | 70.26% | 62.74% | 0.271 | ||
| CAR | 56.36% | 61.80% | 70.31% | 62.82% | 0.279 | ||
| IFD | 57.23% | 61.65% | 69.85% | 62.91% | 0.191 | ||
| perplexity | 57.48% | 63.65% | 70.37% | 63.83% | 0.301 | ||
| Ours | 58.34% | 63.66% | 70.50% | 64.16% | 0.324 | ||
| Ours - Perplexity | All Data | 51.0 | +0.86% | +0.00% | +0.12% | +0.33% | +0.023 |
| Ours - Perplexity | STD > 2.00% | 21.0 | +1.62% | +0.08% | +0.50% | +0.80% | +0.032 |
| Ours - Perplexity | STD > 4.00% | 10.0 | +3.05% | +0.18% | +1.63% | +1.76% | +0.075 |
In this section, we treat each generation strategy from Section 4.3, and response-selection metrics from the related work section, as baselines. We then compare the average training outcomes of our method against these baselines across all tasks.
There are two benchmark sets, detailed in Section 4.2. The first is the Main-experiment corpus, which covers a diverse range of tasks and serves as the primary benchmark for evaluating both the general ranking ability of our metric and the average performance gains achievable by our method.
Since our goal is efficient data selection, we evaluate each metric using only a small subset of the training data. For each method, we repeat the process three times, each time selecting a different subset of size K=50 from the training dataset, and report the average performance across these runs. For example, one run may use the first 50 samples, another the second 50, and so on. The final result is computed as the average of these three evaluations.
@@ 5.1 Hyperparameters
5.2 Evaluation Metrics
Accuracy.
For every {model, dataset} pair, we let each ranking metric select the top-ranked response-generation strategy, fine-tune the model on data produced by that strategy, and record the resulting test accuracy. We then report the macro average of these accuracies across all evaluated tasks. This score answers the practical question: If I trust a metric to choose my training data, how well will my model perform on average?
Weighted Spearman correlation.
To measure how closely a metric’s ranking matches the gold ranking, we compute a weighted Spearman coefficient in which each task is weighted by the standard deviation of accuracies obtained from all candidate strategies; tasks whose choice of strategy matters more thus contribute more. The exact formula and implementation details are provided in Appendix A.4.
5.3 Comparison with Baseline Response Generation Strategies
Table 1 summarizes the average test accuracy obtained when the target model is fine-tuned on data produced by each response-generation strategy. For datasets that provide chain-of-thought (CoT) groundtruth, we additionally evaluate the Rewrite Ground Truth strategy. As this strategy is only applicable to CoT datasets and some datasets do not have CoT groundtruth, its results are excluded from the table to avoid skewing the overall averages; nevertheless, they are included in every metric that ranks candidate strategies on a per-task basis.
Effect of task-specific variance.
Table 13 shows that the performance gap among generation strategies is highly task-dependent: some tasks show differences of several percentage points, while others are nearly insensitive to the chosen strategy. To quantify how much our method helps when the choice of generation strategy matters most, we group every {model, dataset} pair by the standard deviation (SD) of accuracies across baselines. All tasks include all pairs without filtering. High-variance tasks retain only those with . Very-high-variance tasks retain only those with . In the whole Main-experiment corpus, our approach delivers the highest mean accuracy, exceeding the strongest single baseline (Claude) by 0.60% and the mean of all baselines by 1.17%.
When we restrict evaluation to high-variance tasks, the average gain of our method over Claude increases to 1.18%; under the very-high-variance filter, this gain further rises to 2.26%. Compared with the mean of all baselines, the improvements reach 2.41% and 4.54% on the high-variance and very-high-variance subsets, respectively. These results confirm that self-aligned perplexity is particularly valuable in scenarios where candidate generation strategies produce widely divergent training outcomes.
For a more detailed analysis of performance variance across different tasks and models, please refer to Section 6.2.
| Methods | STD Range | num of recorded data | mistral | llama 3 instruct | qwen | Avg Acc | Weighted Spearman Pho |
| Upper bound | All Data | 18.0 | 52.88% | 54.87% | 41.87% | 49.88% | |
| Step-by-step | 37.56% | 44.48% | 31.06% | 37.70% | |||
| GPT-4 ICL examples | 40.86% | 45.84% | 36.86% | 41.19% | |||
| Human examples | 45.01% | 41.82% | 29.89% | 38.91% | |||
| Mini-GPT-4 | 38.68% | 40.29% | 30.51% | 36.49% | |||
| GPT-4 | 39.33% | 41.86% | 31.08% | 37.42% | |||
| Claude | 51.89% | 50.09% | 37.27% | 46.42% | |||
| Ours | 45.29% | 50.06% | 39.66% | 45.00% | |||
| Ours - Claude | -6.60% | -0.03% | +2.39% | -1.42% | |||
| Ours - Avg of Others | +3.07% | +5.99% | +6.89% | +5.31% | |||
| Upper bound | All Data | 18.0 | 52.88% | 54.87% | 41.87% | 49.88% | |
| skywork | 43.66% | 46.01% | 36.12% | 41.93% | 0.101 | ||
| CAR | 41.87% | 44.99% | 35.00% | 40.62% | 0.133 | ||
| IFD | 52.88% | 49.73% | 38.19% | 46.93% | 0.386 | ||
| perplexity | 42.69% | 48.50% | 36.44% | 42.54% | 0.198 | ||
| Ours | 45.29% | 50.06% | 39.66% | 45.00% | 0.241 | ||
| Ours - Perplexity | +2.59% | +1.56% | +3.22% | +2.46% | +0.043 |
5.4 Comparison with Alternative Response Selection Metrics
Table 2 reports results obtained with the same set-up as in Section 5.3, but swapping the ranking metric. Across the full Main-experiment corpus, self-aligned perplexity achieves the best mean accuracy and the highest weighted Spearman correlation; standard perplexity is the closest baseline. All tasks: Using every training run, our metric surpasses standard perplexity by 0.33% in accuracy and by 0.033 in weighted Spearman . High-variance tasks (): The margins widen to 0.80% in accuracy and 0.032 in weighted . Very-high-variance tasks (): Gains further increase to 1.76% in accuracy and 0.075 in weighted . These results mirror the trend observed in Section 5.3: the larger the performance spread among candidate strategies, the more our metric outperforms conventional perplexity, underscoring its value for selecting high-quality training data.
5.5 Performance Differences among Response Generation Strategies Can be Very Large
Candidate response-generation strategies can yield significantly different results depending on the task. To illustrate this, we evaluate various strategies on the PlanBench benchmark introduced in Section 4.2, which is designed to be more difficult than standard instruction-following datasets due to its long-horizon, goal-conditioned reasoning requirements. As shown in Table 3 and Tabel 20(Appendix), training outcomes vary significantly across methods, underscoring the importance of selecting an appropriate generation strategy. Our self-aligned perplexity metric improves accuracy by an average of over standard perplexity and over the mean performance of all strategies. The results further demonstrate that, as the optimal model varies across datasets and continues to shift as APIs evolve, model-aware selection metrics like self-aligned perplexity remain critical.
As shown in Table 3, the average accuracy of IFD is slightly higher than that of our method. However, our method shows comparatively better performance on the main tasks, which cover a broader set of 17 tasks (Table 2). This suggests that our approach may offer more stable and reliable results when applied across a wider range of tasks.
@@ 6.1 Why Self-Aligned Perplexity Outperforms Traditional Perplexity
| Target Response style | Model | Task | PPL | SPPL | SPPL | SPPL |
@@
Traditional perplexity is sensitive to surface-level stylistic cues, so a low score does not necessarily mean the response “feels” familiar to the model. We therefore anchor the metric on the model’s own zero-shot prediction: the closer a candidate lies to this anchor, the more familiar it should be. Injecting that prediction as a single in-context example reshapes the probability distribution, yielding a self-aligned perplexity that more faithfully reflects the response’s true familiarity.
When perplexity fails.
According to Table 4, On ECQA, a deliberately redundant answer(see Appendix A.6 for how we construct this dataset) scores 4.94 in raw perplexity, while the terser, higher-quality GPT-4-direct answer scores 5.55 (Table 4). A similar pattern appears in PIQA (4.55 vs. 6.28). Thus, lower perplexity can sometimes reflect wordiness rather than genuine familiarity with the model’s preferred style.
How self-aligned perplexity helps.
According to Table 4, adding a single in-context example can realign the perplexity scores. For the GPT-4–style response on ECQA, the raw perplexity (PPL) is 5.551, which is higher than the redundant-style response (4.944). After prepending an in-context example drawn from another GPT-4–style answer, the GPT-4 response’s perplexity drops to 4.768. In contrast, when the same example is added to the redundant and step-by-step responses, their perplexities increase from 4.944 and 4.476 to 5.615 and 4.850, respectively.
A similar effect occurs on PIQA: the GPT-4 response on Mistral has an initial perplexity of 6.277, higher than the redundant style (4.570). With a GPT-4 in-context example, its perplexity decreases to 5.962, while the redundant style’s perplexity rises to 6.724.
Across the two tasks, using the model’s own prediction as the in-context anchor consistently lowers the score for its native style by 0.6–1.5 points, restoring the correct ordering and yielding rankings that track downstream fine-tuning gains.
@@
| Ours | Mistral7B | 0.583 | 0.317 |
| TTT (lr=2e-5) | 0.578 | 0.188 | |
| TTT (lr=2e-4) | 0.579 | 0.557 | |
| Ours | Llama3 | 0.637 | 0.288 |
| TTT (lr=2e-5) | 0.613 | 0.168 | |
| TTT (lr=2e-4) | 0.627 | 0.398 | |
| Ours | Qwen2.5 | 0.705 | 0.367 |
| TTT (lr=2e-5) | 0.702 | 0.065 | |
| TTT (lr=2e-4) | 0.709 | 0.273 | |
| Ours | Average | 0.642 | 0.324 |
| TTT (lr=2e-5) | 0.631 | 0.141 | |
| TTT (lr=2e-4) | 0.638 | 0.409 |
| 0–10 | 10–20 | 20–30 | 0–30 | 30–60 | 60–90 | 0–50 | 50–100 | 100–150 | |
| Accuracy (%) | 63.90% | 64.22% | 63.93% | 64.25% | 63.96% | 64.22% | 64.17% | 64.10% | 64.22% |
| Weighted | 0.289 | 0.363 | 0.290 | 0.327 | 0.286 | 0.269 | 0.317 | 0.288 | 0.367 |
| 0–100 | 100–200 | 200–300 | 0–200 | 0–300 | - | - | - | - | |
| Accuracy (%) | 64.18% | 64.09% | 64.08% | 64.13% | 64.13% | - | - | - | - |
| Weighted | 0.332 | 0.330 | 0.323 | 0.333 | 0.342 | - | - | - | - |
6.2 Why Do Some Datasets Show Greater Variance in Training Outcomes?
| Model | gsm8k | math algebra | mmlu | winogrande | piqa | agieval | squad | ecqa | boolq | arc challenge | mmlu pro law | drop | hellaswag | mmlu moral scenarios | math geometry | api bank |
| qwen | 94.8% / 88.9% | 93.4% / 91.6% | 51.0% / 50.9% | 72.7% / 72.1% | 86.0% / 87.9% | 57.2% / 54.5% | 67.3% / 76.2% | 81.6% / 79.3% | 83.6% / 86.5% | 89.3% / 89.0% | 32.1% / 27.6% | 0.7% / 79.4% | 70.6% / 70.6% | 56.8% / 63.3% | 51.9% / 65.8% | 11.7% / 43.3% |
| mistral | 39.3% / 62.5% | 15.4% / 31.9% | 37.6% / 42.2% | 44.9% / 71.3% | 69.2% / 86.9% | 33.5% / 40.0% | 8.8% / 73.2% | 57.6% / 70.0% | 82.3% / 86.7% | 73.5% / 61.1% | 19.4% / 22.9% | 0.6% / 74.6% | 39.7% / 65.4% | 37.2% / 71.3% | 5.1% / 17.7% | 1.8% / 51.0% |
| llama 3 instruct | 85.8% / 81.4% | 58.1% / 56.2% | 43.0% / 48.9% | 68.4% / 69.5% | 81.1% / 86.5% | 43.0% / 43.5% | 24.5% / 75.2% | 72.0% / 72.3% | 78.1% / 88.0% | 79.4% / 80.1% | 25.9% / 27.6% | 15.1% / 79.6% | 56.8% / 72.2% | 33.3% / 67.7% | 18.3% / 27.8% | 0.0% / 49.4% |
We observed a striking regularity across tasks: whenever enlarging the training set from 100 to 1 000 examples yields little or no accuracy gain, the choice of response-generation method matters equally little. Conversely, tasks that continue to improve with more data show pronounced performance gaps between generation strategies.
Let the improvement ratio be defined as , representing the relative gain from increasing the training size ten-fold. Figure 1 plots (x-axis) against the standard deviation of accuracies across generation methods (y-axis). A clear positive trend emerges: once the improvement ratio exceeds roughly 2%, the variance among methods rises sharply; below this threshold, it is nearly zero. We plot Figure 1 using training results from all tasks in the Main-experiment corpus and PlanBench.
The results suggest that divergence across generation strategies is greatest exactly when the dataset still offers headroom for improvement. On such high-variance tasks, selecting the right response-generation method is critical, underscoring the value of our self-aligned perplexity criterion.
Figure 1 shows that many red data points lie to the left of the green 2% threshold line, indicating that training with data generated by different response-generation strategies produces negligible variation in training outcomes on Qwen. We hypothesize that this low-variance pattern arises because the training set contributes little information beyond Qwen’s existing capabilities. This also explains why our method improves accuracy by only 0.45% over the average of all response-generation strategies on Qwen when evaluated on the full dataset. In Table 7, we compare zero-shot accuracy with the accuracy of the three models after being trained on GPT-4o direct-answer data for each task. As shown, only 43.8% of tasks improve after training, suggesting that Qwen already possesses strong prior knowledge of standard benchmarks for most tasks. In contrast, when trained on data with STD > 4%, Qwen achieves a 3.38% accuracy improvement over the average of all other methods. On PlanBench, as demonstrated in Section 5.5 and Table 3, our method yields substantial accuracy gains across the six PlanBench subtasks with Qwen: +3.22% over perplexity and +6.89% over the average of all strategies.
6.3 The effectiveness of Correctness Filter
Correctness filter is helpful mainly on the tasks that the model is very unfamiliar with.
Removing the correctness filter has negligible impact on the 17 main tasks, as shown in the Table 9. The table shows the accuracy and weighted Spearman correlation deltas between the filtered and unfiltered versions of our method (Ours Ours-w/o-Filter); Overall, skipping the filter—i.e., using the model’s raw predictions—only minimally reduces performance in terms of average accuracy and weighted Spearman correlation.
However, removing the correctness filter results in a performance drop on PlanBench, as shown in the table 10.
@@ 6.3 The effectiveness of Correctness Filter
One natural (but computationally expensive) approach to select the optimal response generation strategy is to adopt a Train-Then-Select (TTS) procedure. In this way, we first generate a small dataset (e.g., 100 samples) using each candidate strategy. For each dataset, we train the target model and evaluate its performance. We then rank the strategies based on the results and choose the best-performing one to generate the remainder of the dataset.
When evaluating TTS, we train the target model on 100 samples under two settings: 1) Standard Training: A learning rate of 2e-5 for 20 epochs (matching our main setup). The performance accuracies for each strategy under this setting is in the Table 17. 2) Intense Training: A learning rate of 2e-4 for 40 epochs. The performance accuracies for each strategy under this setting is in Table 18.
After ranking the strategies using TTS, we compare their performance with ours. In Table 5, despite using less data and requiring no training, validation, or testing computations for strategy selection, our method achieves better average accuracy and comparable weighted Spearman correlation.
6.5 Stability of Our Method
As shown in Table 6, accuracy generally improves as the subset size grows, and the overall performance is consistent across ranges. Small subsets sometimes degrade accuracy (values highlighted in red); thus, we recommend using at least 30 samples or even 50 samples for the best performance. Specifically, “0–10”, “10–20”, and “20–30” denote the first, second, and third batches of ten training examples, respectively, while “0–50” and “0–100” correspond to the first 50 and 100 examples. When the subset size reaches 50 or more, average accuracy stabilises. The weighted Spearman correlation () also increases with larger subsets, but the gains taper off once the subset size exceeds 50.